OpenStudy (anonymous):

9x-5x=16 x=4 x=14 x=16 x=40

OpenStudy (some.random.cool.kid):

like terms first

OpenStudy (anonymous):

9-5 is 4 so 4 divided by 16 is ????? that's your answer

OpenStudy (anonymous):

ok so 9x and 5x? @some.random.cool.kid

OpenStudy (some.random.cool.kid):

what @HWBUSTER00

OpenStudy (some.random.cool.kid):

said

OpenStudy (anonymous):

so the like terms are 9x and 5x right? @some.random.cool.kid

OpenStudy (anonymous):

9-5 is 4 so 4x=16 then you divide 4 by 16 @some.random.cool.kid

OpenStudy (some.random.cool.kid):

yes correct and then like @HWBUSTER00 said you minus and then divide

OpenStudy (some.random.cool.kid):

no i wasnt asking lol i was telling her

OpenStudy (anonymous):

okay, so 9-5=4

OpenStudy (anonymous):

then you divide 4 by...?

OpenStudy (anonymous):

16 what is 16/4 .....the answer is x=4 ...........

OpenStudy (anonymous):

ok thnx

OpenStudy (anonymous):

sorry i took so long i had to clean my room real quick